Sierra Company produces its product at a total cost of $89 per unit. Of this amount, $14 per unit is selling and administrative costs. The total variable cost is $58 per unit, and the desired profit is $28 per unit.

Determine the markup percentage using the (a) total cost, (b) product cost, and (c) variable cost methods. Round your answers to one decimal place.

a.  Total cost %
b.  Product cost %
c.  Variable cost %
Answer #1

Mark up percentage On total Cost : 25/89=28.08%

Mark up percentage product Cost:25/(89-14)=33.33%

Mark up percentage variable cost :25/58=43.1%